# Spin-fermion coupling enhances pairing in the pseudogap regime of the hole-doped Hubbard model

### v1

@@ Abstract
With regard to the power law observed when drawing $|\gamma|$ against $-\text{Im}\Sigma$, we note that they are connected via Hedin’s equation $\Sigma\sim GW\gamma$, where $G$ is the Green’s function. Upon a breakdown of Kondo screening, the emergence of poles in the self-energy is ubiquitous in a variety of models of strongly correlated electrons Pavarini et al. (2017). Indeed, such a pole is needed to open a (pseudo)gap without symmetry breaking or order. We observe here that, for the underdoped Hubbard model in two dimensions, $\Sigma$ and $\gamma$ could form poles concomitantly; this raises the intriguing possibility that pairing may be sustained in absence of a Fermi surface, namely, by a Luttinger surface Fabrizio (2022, 2023).
Conclusions.— We computed the spin-fermion coupling $\gamma$ within the 8-site dynamical cluster approximation (DCA) for the hole-doped Hubbard model in the pseudogap regime. We find that $\gamma$ exhibits a significant nodal/antinodal dichotomy in momentum space. The pairing glue is thus enhanced by a large factor $|\gamma|^{2}$ for antinodal fermions, even though these fermions lose their quasiparticle character due to the opening of the pseudogap. Without this enhancement, $d$-wave pairing would not occur in the underdoped regime.
Previous studies of the spin-fluctuation glue in the normal state often assumed an isotropic spin-fermion coupling $\gamma\approx 1$ (e.g., Maier et al. (2007a, b); Mishra et al. (2014); Restrepo et al. (2022)), which we find may only apply in weakly correlated systems or at high temperatures. With regard to the superconducting state, a recent study found that spin fluctuations contribute only about half of the superconductivity in the strongly correlated regime, based on an isotropic $\gamma$ Dong et al. (2022b). It is plausible that for underdoping the dichotomy and antinodal enhancement of $\gamma$ observed here persist in the superconducting state  [^1] Note that spin fluctuations may change qualitatively in the superconducting state Inosov et al. (2009)., calling for further investigation of the lightly doped Hubbard model.

### v2

@@ Abstract
With regard to the power law observed when drawing $|\gamma|$ against $-\text{Im}\Sigma$, we note that they are connected via Hedin’s equation $\Sigma\sim GW\gamma$, where $G$ is the Green’s function. Upon a breakdown of Kondo screening, the emergence of poles in the self-energy is ubiquitous in a variety of models of strongly correlated electrons Pavarini et al. (2017). Indeed, such a pole is needed to open a (pseudo)gap without symmetry breaking or order. We observe here that, for the underdoped Hubbard model in two dimensions, $\Sigma$ and $\gamma$ could form poles concomitantly; this raises the intriguing possibility that pairing may be sustained in the absence of a Fermi surface, namely, by a Luttinger surface Fabrizio (2022, 2023); Worm et al. (2024).
Conclusions.— We computed the spin-fermion coupling $\gamma$ within the eight-site dynamical cluster approximation (DCA) for the hole-doped Hubbard model in the pseudogap regime. We find that $\gamma$ exhibits a significant nodal/antinodal dichotomy in momentum space. The pairing glue is thus enhanced by a large factor $|\gamma|^{2}$ for antinodal fermions, even though these fermions lose their quasiparticle character due to the opening of the pseudogap. Without this enhancement, $d$-wave pairing would not occur in the underdoped regime.
Previous studies of the spin-fluctuation glue in the normal state often assumed an isotropic spin-fermion coupling $\gamma\approx 1$ (e.g., Refs. Maier et al. (2007a, b); Mishra et al. (2014); Restrepo et al. (2022)), which we find may only apply in weakly correlated systems or at high temperatures. With regard to the superconducting state, a recent study found that spin fluctuations contribute only about half of the superconductivity in the strongly correlated regime, based on an isotropic $\gamma$ Dong et al. (2022b). It is plausible that for underdoping the dichotomy and antinodal enhancement of $\gamma$ observed here persist in the superconducting state  [^2] Note that spin fluctuations may change qualitatively in the superconducting state Inosov et al. (2009)., calling for further investigation of the lightly doped Hubbard model.